OpenStudy (anonymous):

-7+3x+4x Simplify :)

OpenStudy (gabebae):

Add like terms?

pooja195 (pooja195):

Combine like terms 3x+4x=?

TheSmartOne (thesmartone):

Hint: 3x + 4x = (3+4)x = ??

OpenStudy (anonymous):

what about the -7?

pooja195 (pooja195):

It remains the same.

OpenStudy (anonymous):

ok thanks

TheSmartOne (thesmartone):

Anytime :D